Tohban Report 2013-03-27

From RHESSI Wiki

Jump to: navigation, search

Tohban Reports
Start Date: 20 Mar 2013
End Date: 27 Mar 2013
Tohban: Lindsay Glesener
Tohban email:
Next Tohban: Mitsuo Oka
List all reports


Solar Activity

It was another quiet week. There are four active regions on the disk and nothing major about to rotate on. A few days ago there was a flurry of C flares, but that AR quickly rotated off. A new AR about to come on-disk is produced an occulted B flare, but quiet activity is expected to continue.

How many GOES flares occurred?

 Flares above B, C, M, X class were     24    19     1     0

Memory Management

With one exception (see D5 update below) the SSR emptied after each pass set.

Spacecraft Status

The spin rate has dropped to ~14.6 rpm, so a spinup will be performed this afternoon (nominally to 15 rpm).

Data Gaps

No gaps!

Detector issues

At 01:20 UT on March 25, there was an arc in detector 5 that set off an interval of noisy triggers. D5 front fast and slow rates jumped up to the max and stayed there for ~1.5 hours before starting to settle. Over the next few hours the rates gradually settled back to a steady-state elevated, but not "dangerous" level. Currently D5 rates are higher than the other detectors but livetime is only being degraded by <1%. Note that the detector settled on its own; although it was closely watched no corrective actions were taken the past few days. At the Ops meeting today we decided to lower the HV for D5 by ~100V as a token gesture to prevent future arcs. For now, we will leave the thresholds as they are.

A similar event happened in November of 2011 (which was before the last anneal); see the tohban reports for more info. As with that event, the excess noise here is at low energies, as you can see in the first plot below.

Browser link to the time of the arc: [1]

Some plots illustrating the issue:

2013mar25-D5-data.png 2013mar25-D5-fast.png 2013mar25-D5-slow.png

Other notes

The data processing chain generating quicklook info has been hung up since the 19th. Jim will fix this and will also use this opportunity to update the flare list. This issue affects the quicklook time profiles in Browser, but the monitor rates and actual event data (i.e. for data analysis) are all present.

Decimation Normal/vigorous
HLAT Decimation Rear decimation weight 6, no front decimation
Night time data (fronts) +/- 4 minutes
Night time data (rears) Taking data at nights
Require extra passes? No
Requirement for moving pointer? No
Attenuator operation Normal
Detector problems? Detector 5 arc + extended noise period.
Personal tools